# The Impact of Differential Feature Under-reporting on Algorithmic Fairness

### v1

@@ 5. Impact on selection rate disparity
We study the effect of differential feature under-reporting on selection rate disparities in linear regression. For the remainder of this section, we assume that $Z\perp G$. Similar to our previous discussion, we assume a $d$-dimensional feature setting in which only the first feature is subject to under-reporting while all other features are fully observed. We further assume that features are jointly Gaussian, i.e. $Z\sim\mathcal{N}\left(\mu,\Sigma\right)$ where $\mu\in\mathds{R}^{d}$ and the covariance matrix $\Sigma\in\mathds{R}^{d\times d}$ is positive definite. This has the benefit that predictions $\hat{Y}=\hat{Y}_{X}$ follow a Gaussian mixture distribution which allows us to directly analyze group selection rates. If under-reporting rates are the same across groups, i.e. $m_{1}^{0}=m_{1}^{1}$, there is no selection rate disparity as both groups have the same distributions of features and predictions. If the under-reporting rates vary between groups, we observe the following.
Proposition 1 ().
@@
We refer to Proposition A.2 for an expanded version of this finding including a discussion of sufficiently high thresholds. Proposition 1 shows that over-selection primarily depends on the variance in predictions. When cutting off at a high threshold, the group with more feature under-reporting is over-selected if the variance in predictions for examples with defaulted feature exceeds the prediction variance for fully observed examples (Case 1). It is under-selected if the variance of predictions is larger for the examples with fully observed features (Case 2). Intuitively, we may expect that, in prediction settings with high thresholds, information deficiency in a group always leads to under-selection of that group because it prompts the group’s risk distribution to concentrate more closely around its mean moving more mass below the threshold. We find that, analytically, outcome disparities can go into either direction and sometimes groups with information deficiency are over-selected. While our findings suggest that this is mostly a question of variance in predictions, this is likely only part of the story in more general cases with group-dependent feature distributions. More general settings are studied empirically in Section 7.
@@
[…]
@@
Thresholds are considered high if they exceed the turning point defined in Proposition A.2. The corollary shows that over-selection due to feature under-reporting depends on the signs and magnitudes of the true parameters $\beta$ and the covariances between features. If $\text{sign}\left(\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]\right)=1,$ If the sign is negative, the group with more under-reporting is over-selected if the covariance-weighted sum of true parameters is sufficiently large in absolute value. Otherwise, feature under-reporting still leads to under-selection.
@@ 6. Solution approaches
Sections 4 and 5 show how ignoring differential feature under-reporting can lead to unfair selection rates across groups. In the following, we explore how conventional missing data methods can be adapted to the under-reporting setting. We then propose a new set of methods that is specifically tailored to this setting by separating the problem into two steps — estimation and prediction. For the estimation step, we provide a method that recovers the ground truth data generating model from observed data. For the prediction step, we derive optimal group-dependent imputation values. As before, we assume under-reporting occurs only in the first feature which is observed as $X_{1}=Z_{1}\xi_{1}$.
@@
    Feature omission. Discarding mismeasured feature vector $X_{1}$ entirely mitigates the bias introduced by under-reporting without requiring missingness indicators. While this may be a viable option in some settings, it can lower model performance significantly and the fairness impact is non-obvious. When assuming a linear ground truth, feature omission leads to omitted variable bias in parameter estimates $\hat{\beta}_{[2:d]}$ which has been studied previously in the econometrics literature (Angrist and Pischke, 2008) (also see Appendix B).
@@
The fact that the augmented loss is unbiased with respect to under-reporting noise implies that a prediction model on observed data estimated with augmented loss, i.e. $\hat{f}=\arg\min_{f\in\mathcal{F}}\mathds{E}_{(X,Y)}\left[\tilde{l}(f,X,Y)\right],$ asymptotically recovers the Bayes optimal model on the true features $Z$. If $Y=\alpha+\beta^{T}Z$, $\mathcal{F}$ is the class of linear functions $f:\mathds{R}^{d}\to\mathds{R}$ and $l(f,z,y)=(f(z)-y)^{2}$ denotes squared error loss, the true parameters $\hat{\alpha}=\alpha$ and $\hat{\beta}=\beta$ are retrieved. Note that squared error loss is not bounded and estimating $\hat{f}$ requires the additional constraint $\tilde{l}(f,X,Y)\geq 0$. Lemma 1 operates in a group-agnostic setting with $Z\bot G$. Lemma A.3 provides a more general group-dependent version of the finding.
@@
which can be estimated directly from observed data if the under-reporting rate $1-m$ is known. If feature distributions vary across groups, group-dependent optimal prediction imputation values can be derived as described in Lemma A.4.
@@
Both augmented loss estimation and optimal prediction imputation require access to the under-reporting rate $1-m$ which is typically unknown. In some cases, it may be possible to obtain supplementary data that can be used to estimate $m$. For example, in administrative data with under-reported health features for privately insured individuals, an external private insurance health claims dataset could be used to estimate the expected rate of true 0s. In most settings, estimation of under-reporting rates needs to rely directly on the observed data. Assume we have access to a dataset $V=\{(x,y)_{i=1}^{n}\}$. We split $V$ into a training portion $V_{\text{train}}$ and evaluation portion $V_{\text{eval}}$. Let $P$ denote the subset of examples from $V_{\text{eval}}$ for which $x_{1}\neq 0$. We draw on the literature on Positive and Unlabeled (PU) learning (Elkan and Noto, 2008) and estimate under-reporting rates as follows. First, we fit a model $h$ on $V_{\text{train}}$ to estimate $P(X_{1}\neq 0\mid X_{[2:d]},Y)$. Second, we evaluate $h$ on $P_{\text{eval}}$. The estimator for the share of observed values $m$ is given by $\hat{m}=\frac{1}{\mid P_{\text{eval}}\mid}\sum_{(x,y)\in P_{\text{eval}}}h(x_{[2:d],y}).$ The estimation procedure assumes that under-reporting occurs completely at random. Our experimental setting assumes under-reporting completely at random within groups and thus $m_{0}$ and $m_{1}$ can be estimated with the described procedure by restricting $V$ to examples from the respective group.
@@ 7.1. Publicly available datasets
Both COMPAS data (Angwin et al., 2016) and German credit data (Repository, 1994) are widely used across the algorithmic fairness literature. The American Community Survey (ACS) Income dataset is comprised of 2018 census data from California queried using the folktables package (Ding et al., 2021). The datasets vary in size, number of features, and prediction tasks as shown in Table D.1. We conduct all experiments with both gender and race as group columns if available. All models are group-blind and race information is never included as predictive feature. While the results for ACS Income and COMPAS are covered in the main text, we refer to Appendix F.1 for a discussion of the German credit data experiments.
@@
Since all of the prediction tasks are binary classification, we opt to generate semi-synthetic regression labels for our experiments. We first fit a logistic regression model to the entire dataset and extract the fitted probabilities. For the ACS Income data, the probabilities are rescaled to more closely reflect the \$50,000 income threshold. Next, we fit a linear regression model using the same features and the predicted probabilities as outputs. The fitted values from this linear model are chosen as the new “true” labels for our experiment. This outcome augmentation procedure allows us to generate artificial settings with a truly linear ground truth similar to the settings studied in Sections 4 and 5 while leaving realistic covariance structures intact. We further experiment with controlling the $R^{2}$ of the true linear model by adding additional noise to outcomes and report fairness implications in Appendix F.2.
@@
Regression models are trained to predict semi-synthetic outcomes based on the features of the respective dataset. We then select the top $C$ share of the predictions as high risk and evaluate excess selection rates as given in Definition 1 to assess the fairness impact of under-reporting. Artificial under-reporting is added to one feature column at a time and we repeat the experiments for each outcome column, group column, and under-reporting rate. Under-reporting rates range from 0-90% in 10 percentage point increments, and we add under-reporting to only one group at a time (e.g., we set 10% of a feature in the male group to 0 while leaving the features of the female group unchanged). Only numeric features are considered for under-reporting since, in administrative data, binary features are often categorical dummies or thresholded versions of continuous count features. All models are trained with 80% of the datasets while withholding 20% for testing. In addition to the baseline models, we experiment with various solution approaches as described in Section 6. This includes our proposed methods of group-dependent augmented loss estimation and group-dependent optimal prediction imputation.
@@ 7.2. County-level birth data
We present an analysis of a private administrative dataset we obtained from a county in the US. The dataset contains information on newborn children and their families including demographics, child protective services history, birth record data, and mental and behavioral health information for those who used publicly funded services. We set up a prediction task that attempts to mirror the analysis described in the Hello Baby model methodology report from Allegheny County (for Social Data Analytics at the Auckland University of Technology, 2020). The Hello Baby model was developed to predict which families are at greatest risk of having their child removed by Child Protective Services (CPS) during their first three years of life, and is used to prioritize families for opt-in, voluntary supportive services. Using our data, we train a similar model, and explore the effect of adding additional under-reporting to the behavioral and mental health data fields.
@@
We use the birth dataset with its original prediction outcomes to showcase a realistic example of the effect of under-reporting. As before, the data is separated into 80% for training and 20% for testing. We fit separate logistic regression models on three datasets. (1) Data as observed. (2) Data with behavioral health features set to 0 for privately insured individuals, i.e. mothers that are not insured through Medicaid. (3) Data without behavioral health features. For illustration, results are stratified by whether individuals are covered by Medicaid, and by whether individuals identify as African-American. Medicaid coverage and race are not used as features in any of the models.
@@ 8. Results
Figure 2 summarizes the results for the experiments on the ACS Income data. We see that feature under-reporting in ‘education attainment’ and ‘hours worked per week’ consistently leads to under-selection of the group with under-reporting. This is true irrespective of whether feature under-reporting is injected into the female sub-group or the male sub-group, and we observe the same effect when under-reporting is added based on the individuals’ racial group. The figure additionally suggests that more under-reporting generally leads to increasing under-selection. Intuitively, it makes sense that both education attainment and hours worked per week contribute positively to predicted income which is confirmed by the parameter estimates (Figure D.3). Exploration of the covariance matrix of the unbiased features further reveals that all numeric columns in the dataset are positively correlated which together creates a setting reminiscent of the Case 2 scenario studied in Section 5. On a high-level, the scenario predicts that the group with more under-reporting is under-selected in the given setting which aligns with our observations. In addition to selection rate disparity, feature under-reporting in the data also leads to decreased model accuracy as displayed in Figure D.2, and the parameter estimates in Figure D.3 display an attenuation effect as predicted in Section 4.
@@
We focus on results for under-reporting in count features (Figure D.1) and point to Appendix F.3 for additional results. The feature ‘priors count’, i.e. the number of previous criminal offenses individuals have been convicted of, emerges as important feature with respect to under-reporting. Under-reporting in priors count leads to under-selection of the impacted group. This pattern repeats itself for any of the groups and both of the available prediction outcomes. The more feature under-reporting in a group, the larger the occurring outcome disparity. Similarly to the previous results, this suggests a setting of Case 2 as discussed in Section 5. As before, parameter estimates suggest an attenuation effect which is displayed in Figure 3. Under-reporting in priors count could be interpreted as an extreme case of crimes that do not result in arrest. Assuming that one demographic group is more likely to be convicted for committed crimes than the other group, the result implies that the already more frequently targeted group may additionally be flagged as high risk for recidivism at disproportionate rates. Racial disparities in arrest rates and police encounters are well-documented in the US (e.g. Alexander, 2010; Butcher et al., 2022; Fogliato et al., 2021; Pierson et al., 2020) which highlights the importance of this finding.
@@
Our experiments reveal that none of the standard missing data methods reliably mitigate the bias introduced through under-reporting and, instead, can have significant effects on excess selection rates themselves. Omission of the feature ‘h worked per week’ leads to over-selection of the female group and under-selection of the male group in the ACS Income data. This is because female individuals report to work on average less than male individuals (35.43h/week vs. 40.05h/week) while work hours contribute positively to income (Figure D.3). Omitting the feature blinds the model to these differences. Similar effects occur with the feature ‘education attainment’, and ‘priors count’ in the COMPAS data (Figures D.1 and 3).
[…]
@@
Figure 3. Excess selection rates of group Other (i.e. not African-American) (left columns), parameter estimates (middle column), and test set $R^{2}$ (right columns) when under-reporting is injected into ‘priors count’ in group Other using the COMPAS dataset and synthetic two-year recidivism outcomes. In (a), the model is trained and evaluated using the under-reported feature. For (b), we first train a multiple imputation model and then train and evaluate the prediction model using probabilistic imputations. For (c), the model is trained on only rows without 0-entries in ‘priors count’ and evaluated on the under-reported data. In (d), we train with group-dependent augmented loss and use group-dependent optimal imputation values for prediction. Results are reported as averages over 30 runs. Shaded areas correspond to one standard deviation. The solid dots in the middle column correspond to true parameters. Note that in order to preserve readability, parameter estimates are only displayed for continuous features. Figure D.6 provides an overlay plot of the rightmost column for easy comparison.
@@
For multiple imputation on the COMPAS data, we see that the excess selection rate flips signs and the group with under-reported ‘priors count’ is over-selected (Figures 3 and D.4). This is because the feature has a lot of true 0-entries that are wrongfully imputed as positive values. The cost incurred by these wrong imputations exceeds the benefit of imputation. In comparison to training on mismeasured features directly, the parameter estimation bias is considerable even for small amounts of under-reporting. For high under-reporting rates, the excess selection rate follows a similar pattern as the excess selection rate with feature omission since imputation is conducted using the features already present in the model adding little to no additional information. Since our models are well-specified, row omission recovers the true parameter estimates as displayed for the COMPAS data and feature ‘priors count’ in Figure 3. Despite access to the ground-truth, we observe that under-reporting bias introduced at prediction time increases the selection rate disparity. While the model without row omission is able to shift weight to correlated features as more and more entries for ‘priors count’ are under-reported, the row omission model cannot make use of the feature correlations ultimately leading to the increasing rather than decreasing disparities. With the same reasoning, the test set performance as measured by $R^{2}$ is decreased as displayed in the figure.
@@
We contrast the performance of our method and standard approaches for handling missing features at the example of the ‘priors count’ feature in the COMPAS dataset. Under-reporting rates are estimated with the procedure described in Section 6. We refer to Appendix E for further details on the under-reporting rate estimation. The results in Figure 3 show that selection rate disparities decrease considerably when using group-dependent augmented loss and group-dependent optimal prediction imputation. In contrast multiple imputation, this fairness improvement comes at no visible cost in performance. In fact, the average test set $R^{2}$ of the corrected model is very similar to, and even slightly higher than, the test set $R^{2}$ of the model trained directly on under-reported data (see Figure D.6). Despite some variability, the average parameter estimates of the corrected model appear more stable across different amounts of under-reporting which suggests that the method successfully diminishes the bias introduced by under-reporting.
@@
For the birth data, Figure 4 suggest that under-reporting of all behavioral health data for the non-Medicaid population leads to over-selection of the Medicaid population. In the displayed overall selection rate range (¡10%), the Medicaid population is selected about 10% more often than in the “true” data setting. Note that in reality this difference could be even larger because some of the “true” data features were likely already under-reported. As shown in the Figure, some of the resulting disadvantage is still observable when evaluating performance in the group of Black individuals. This can be explained by the fact that the two group variables are positively correlated in the dataset ($\rho=0.44$). Removing the behavioral health columns altogether leads to a reversal of selection rate disparities in the Medicaid / non-Medicaid groups, and significantly increases selection of the Black sub-population. This gives further evidence that feature omission is not a successful remedy for the disparities caused by differential under-reporting.

### v2

@@ 5. Impact on selection rate disparity
We study the effect of differential feature under-reporting on selection rate disparities in linear regression. Similar to our previous discussion, we assume a $d$-dimensional feature setting in which only the first feature is subject to under-reporting and $Z\perp G$. We further assume that features are jointly Gaussian, i.e. $Z\sim\mathcal{N}\left(\mu,\Sigma\right)$ where $\mu\in\mathds{R}^{d}$ and the covariance matrix $\Sigma\in\mathds{R}^{d\times d}$ is positive definite. This has the benefit that predictions $\hat{Y}=\hat{Y}_{X}$ follow a Gaussian mixture distribution which allows us to directly analyze group selection rates. If under-reporting rates are the same across groups, there is no selection rate disparity as both groups have the same feature distributions. If the under-reporting rates vary between groups, we observe the following.
Proposition 1.
@@
We refer to Proposition B.2 for an expanded version of this finding including a discussion of sufficiently high thresholds. Proposition 1 shows that over-selection primarily depends on variance in predictions. When cutting off at a high threshold, the group with more feature under-reporting is over-selected if the variance in predictions for examples with defaulted feature exceeds the prediction variance for fully observed examples (Case 1). It is under-selected if the variance of predictions is larger for the examples with fully observed features (Case 2). Intuitively, at high thresholds, information deficiency in a group always leads to under-selection because it prompts the group’s risk distribution to concentrate more closely around its mean moving more mass below the threshold. We find that, analytically, outcome disparities can go into either direction and sometimes groups with information deficiency are over-selected. While our findings suggest that this is mostly a question of variance in predictions, this is likely only part of the story in settings with group-dependent feature distributions. We study more general settings empirically in Section 7.
@@
$\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]<-c,$
@@
Thresholds are considered high if they exceed the turning point defined in Proposition B.2. The corollary shows that over-selection due to feature under-reporting depends on the signs and magnitudes of the true parameters $\beta$ and the covariances between features. If $\text{sign}\left(\frac{1}{\beta_{1}}\sum_{j=2}^{d}\beta_{j}\text{Cov}\left[{Z_{1}},{Z_{j}}\right]\right)=1,$ e.g. if all true parameters and covariances are non-negative, the group with more missingness will always be under-selected at high thresholds. If the sign is negative, the group with more under-reporting is over-selected if the covariance-weighted sum of true parameters is sufficiently large in absolute value. Otherwise, feature under-reporting still leads to under-selection.
@@ 6. Solution approaches
Sections 4 and 5 show how ignoring differential feature under-reporting can lead to disparities in selection rates across groups. In the following, we explore how conventional missing data methods can be adapted to the under-reporting setting. We then propose a new set of methods that is specifically tailored to this setting by separating the problem into two steps—estimation and prediction. For the estimation step, we provide a method that recovers the ground truth data generating model from observed data. For the prediction step, we derive optimal group-dependent imputation values. As before, we assume under-reporting occurs only in the first feature which is observed as $X_{1}=Z_{1}\xi_{1}$.
@@
    Feature omission. Discarding the mismeasured feature vector $X_{1}$ doesn’t require missingness indicators and mitigates the bias introduced by under-reporting. However, this approach may decrease model performance significantly, and may itself introduce bias. When assuming a linear ground truth, feature omission leads to omitted variable bias in parameter estimates $\hat{\beta}_{[2:d]}$, which has been studied previously in the econometrics literature (Angrist and Pischke, 2008) (also see Appendix C).
@@
The fact that the augmented loss is unbiased with respect to under-reporting noise implies that a prediction model on observed data estimated with augmented loss, i.e. $\hat{f}=\arg\min_{f\in\mathcal{F}}\mathds{E}_{(X,Y)}\left[\tilde{l}(f,X,Y)\right],$ asymptotically recovers the Bayes optimal model on the true features $Z$. If $Y=\alpha+\beta^{T}Z$, $\mathcal{F}$ is the class of linear functions $f:\mathds{R}^{d}\to\mathds{R}$ and $l(f,z,y)=(f(z)-y)^{2}$ denotes squared error loss, the true parameters $\hat{\alpha}=\alpha$ and $\hat{\beta}=\beta$ are retrieved. Note that squared error loss is not bounded and estimating $\hat{f}$ requires the additional constraint $\tilde{l}(f,X,Y)\geq 0$. Lemma 1 operates in a group-agnostic setting with $Z\bot G$. Lemma B.3 provides a more general group-dependent version of the finding.
@@
which can be estimated directly from observed data if the under-reporting rate $1-m$ is known. If feature distributions vary across groups, group-dependent optimal prediction imputation values can be derived as described in Lemma B.4.
@@
Both augmented loss estimation and optimal prediction imputation require access to the reporting rate, $m$, which is typically unknown. In some cases, it may be possible to obtain supplementary data that can be used to estimate $m$. For example, in administrative data with under-reported health features for privately insured individuals, an external private insurance health claims dataset could be used to estimate the expected rate of true 0’s. In most settings, estimation of under-reporting rates needs to rely directly on the observed data. Assume we have access to a dataset $V=\{(x,y)_{i=1}^{n}\}$. We split $V$ into a training portion $V_{\text{train}}$ and evaluation portion $V_{\text{eval}}$. Let $P_{\text{eval}}$ denote the subset of examples from $V_{\text{eval}}$ for which $x_{1}\neq 0$. We draw on the literature on Positive and Unlabeled (PU) learning (Elkan and Noto, 2008) and estimate under-reporting rates as follows. First, we fit a model $h$ on $V_{\text{train}}$ to estimate $P(X_{1}\neq 0\mid X_{[2:d]},Y)$. Second, we evaluate $h$ on $P_{\text{eval}}$. The estimator for the share of observed values $m$ is given by $\hat{m}=\frac{1}{\mid P_{\text{eval}}\mid}\sum_{(x,y)\in P_{\text{eval}}}h(x_{[2:d]},y).$ The estimation procedure assumes that under-reporting occurs completely at random. Our experimental setting assumes under-reporting completely at random within groups and thus $m_{0}$ and $m_{1}$ can be estimated with the described procedure by restricting $V$ to examples from the respective group. For more details on the estimation procedure, we refer to Appendix F.
@@ 7.1. Publicly available datasets
Both COMPAS data (Angwin et al., 2016) and German credit data (Repository, 1994) are widely used across the algorithmic fairness literature. The American Community Survey (ACS) Income dataset is comprised of 2018 census data from California (Ding et al., 2021). Datasets vary in size, number of features, and prediction tasks as shown in Table E.2. We conduct all experiments with both gender and race as group columns if available. All models are group-blind and race information is never included as predictive feature. Results for the German credit dataset are discussed in Appendix G.1.
@@
Since all of the prediction tasks are binary classification, we opt to generate semi-synthetic regression labels for our experiments. We first fit a logistic regression model to the entire dataset and extract the fitted probabilities. For the ACS Income data, the values are rescaled to center around the \$50,000 income threshold. Next, we fit a linear regression model using the same features and the predicted probabilities as outputs. The fitted values from this linear model are chosen as the new “true” labels for our experiment. This outcome augmentation procedure allows us to generate artificial settings with a truly linear ground truth similar to the settings studied in Sections 4 and 5 while leaving realistic covariance structures intact. We further experiment with controlling the $R^{2}$ of the true linear model by adding additional noise to outcomes and report fairness implications in Appendix G.2.
@@
Regression models are trained to predict semi-synthetic outcomes based on the features of the respective dataset. We select the top $C$ share of the predictions as high risk and evaluate excess selection rates to assess the fairness impact of under-reporting. Artificial under-reporting is added to one feature column at a time and we repeat the experiments for each outcome column, group column, and under-reporting rate. Under-reporting rates range from 0-90% in 10 percentage point increments, and we add under-reporting to only one group at a time (e.g., we set 10% of a feature in the male group to 0 while leaving the features of the female group unchanged). Only numeric features are considered for under-reporting since, in administrative data, binary features are often categorical dummies or thresholded versions of continuous count features. All models are trained with 80% of the datasets while withholding 20% for testing. We experiment with various solution approaches as described in Section 6. This includes our proposed methods of group-dependent augmented loss estimation and group-dependent optimal prediction imputation.
@@ 7.2. County-level birth data
We present an analysis of a private administrative dataset we obtained from a county in the US. The dataset contains information on newborn children and their families including demographics, child protective services history, birth record data, and mental and behavioral health information for those who used publicly funded services. We set up a prediction task that attempts to mimic the analysis described in the Hello Baby model methodology report from Allegheny County (hello-baby). The Hello Baby model was developed to predict which families are at greatest risk of having their child removed by Child Protective Services (CPS) during their first three years of life, and is used to prioritize families with newborn children for opt-in, voluntary supportive services. Using our data, we train a similar model, and explore the effect of adding additional under-reporting to the behavioral and mental health data fields.
@@
We use the birth dataset with its original prediction outcomes to showcase a realistic example of the effect of under-reporting. As before, the data is separated into 80% for training and 20% for testing. We fit separate logistic regression models on three datasets. (1) Data as observed. (2) Data with behavioral health features set to 0 for privately insured individuals, i.e. mothers that are not insured through Medicaid. (3) Data without behavioral health features. For illustration, results are stratified by whether individuals are covered by Medicaid, and by whether the mother’s race is recorded as Black. Medicaid coverage and race are not used as features in any of the models.
@@ 8. Results
Figure 2 summarizes the results for the experiments on the ACS Income data. We see that feature under-reporting in ‘education attainment’ and ‘hours worked per week’ consistently leads to under-selection of the group with under-reporting. This is true irrespective of whether feature under-reporting is injected into the female sub-group or the male sub-group, and we observe the same effect when under-reporting is added based on the individuals’ racial group. The figure additionally suggests that more under-reporting generally leads to increasing under-selection. Intuitively, it makes sense that both education attainment and hours worked per week contribute positively to predicted income which is confirmed by the parameter estimates (Figure E.3). Exploration of the covariance matrix of the unbiased features further reveals that all numeric columns in the dataset are positively correlated which together creates a setting reminiscent of the Case 2 scenario studied in Section 5. At a high-level, our theoretical analysis predicts that the group with more under-reporting is under-selected in this setting which aligns with our observations. In addition to selection rate disparity, feature under-reporting in the data also leads to decreased model accuracy as displayed in Figure E.2, and the parameter estimates in Figure E.3 display an attenuation effect as predicted in Section 4.
@@
We focus on results for under-reporting in count features (Figure E.1) and point to Appendix G.3 for additional results. The feature ‘priors count’, i.e. the number of previous criminal offenses individuals have been convicted of, emerges as important feature with respect to under-reporting. Under-reporting in priors count leads to under-selection of the impacted group. This pattern repeats itself for any of the groups and both of the available prediction outcomes. The more feature under-reporting in a group, the larger the occurring outcome disparity. Similarly to the previous results, this suggests a setting of Case 2 as discussed in Section 5. As before, parameter estimates suggest an attenuation effect which is displayed in Figure 3. Under-reporting in priors count could be interpreted as an extreme case of crimes that do not result in arrest. Assuming that one demographic group is more likely to be convicted for committed crimes than the other group, the result implies that the already more frequently targeted group may additionally be flagged as high risk for recidivism at disproportionate rates. Racial disparities in arrest rates and police encounters are well-documented in the US (e.g. Alexander, 2010; Butcher et al., 2022; Fogliato et al., 2021; Pierson et al., 2020) which highlights the importance of this finding.
@@
Our experiments reveal that none of the standard missing data methods reliably mitigate the bias introduced through under-reporting and, instead, may themselves introduce disparities in selection rates. Omission of the feature ‘hours worked per week’ leads to over-selection of the female group and under-selection of the male group in the ACS Income data. This is because female individuals report to work on average less than male individuals (35.43h/week vs. 40.05h/week) while work hours contribute positively to income (Figure E.3). Omitting the feature blinds the model to these differences. Similar effects occur with the feature ‘education attainment’, and ‘priors count’ in the COMPAS data (Figures E.1 and 3).
(a) Estimation and prediction with under-reported feature
(b) Estimation and prediction with multiple imputation
(c) Estimation on rows with non-zero entries and prediction with under-reported feature
(d) Estimation with group-dependent augmented loss and prediction with optimal group-dependent imputation (our method)
@@
Figure 3. Excess selection rates of group Other (i.e. not African-American) (left columns), parameter estimates (middle column), and test set $R^{2}$ (right columns) when under-reporting is injected into ‘priors count’ in group Other using the COMPAS dataset and synthetic two-year recidivism outcomes. In (a), the model is trained and evaluated using the under-reported feature. For (b), we first train a multiple imputation model and then train and evaluate the prediction model using probabilistic imputations. For (c), the model is trained on only rows without 0-entries in ‘priors count’ and evaluated on the under-reported data. In (d), we train with group-dependent augmented loss and use group-dependent optimal imputation values for prediction. Results are reported as averages over 30 runs. Shaded areas correspond to one standard deviation. The solid dots in the middle column correspond to true parameters. Note that in order to preserve readability, parameter estimates are only displayed for continuous features. Figure E.6 provides an overlay plot of the rightmost column for easy comparison.
@@
For multiple imputation on the COMPAS data, we see that the excess selection rate flips signs and the group with under-reported ‘priors count’ is over-selected (Figures 3 and E.4). This is because the feature has a lot of true 0-entries that are wrongfully imputed as positive values. The cost incurred by these wrong imputations exceeds the benefit of imputation. In comparison to training on mismeasured features directly, the parameter estimation bias is considerable even for small amounts of under-reporting. For high under-reporting rates, the excess selection rate follows a similar pattern as the excess selection rate with feature omission since imputation is conducted using the features already present in the model adding little to no additional information. Since our models are well-specified, row omission recovers the true parameter estimates as displayed for the COMPAS data and feature ‘priors count’ in Figure 3. Despite access to the ground-truth, we observe that under-reporting bias introduced at prediction time increases the selection rate disparity. While the model without row omission is able to shift weight to correlated features as more and more entries for ‘priors count’ are under-reported, the row omission model cannot make use of the feature correlations ultimately leading to the increasing rather than decreasing disparities. With the same reasoning, the test set performance as measured by $R^{2}$ is decreased as displayed in the figure.
@@
We contrast the performance of our method and standard approaches for handling missing features at the example of the ‘priors count’ feature in the COMPAS dataset. Under-reporting rates are estimated with the procedure described in Section 6. We refer to Appendix F for further details on the under-reporting rate estimation. The results in Figure 3 show that selection rate disparities decrease considerably when using group-dependent augmented loss and group-dependent optimal prediction imputation. In contrast multiple imputation, this fairness improvement comes at no visible cost in performance. In fact, the average test set $R^{2}$ of the corrected model is very similar to, and even slightly higher than, the test set $R^{2}$ of the model trained directly on under-reported data (see Figure E.6). Despite some variability, the average parameter estimates of the corrected model appear more stable across different amounts of under-reporting which suggests that the method successfully diminishes the bias introduced by under-reporting.
@@
For the birth data, Figure 4 suggest that under-reporting of all behavioral health data for the non-Medicaid population leads to over-selection of the Medicaid population. In the displayed overall selection rate range (¡10%), the Medicaid population is selected about 10% more often than in the “true” data setting. Note that in reality this difference could be even larger because some of the “true” data features were likely already under-reported. As shown in the Figure, some of the resulting disadvantage is still observable when evaluating performance for Black families. This can be explained by the fact that the two group variables are positively correlated in the dataset ($\rho=0.44$). Excluding behavioral health features altogether leads to a reversal of selection rate disparities in the Medicaid / non-Medicaid groups, and significantly increases selection of the Black sub-population for overall selection rates less than about 7%. This provides additional evidence highlighting that omitting features is an unreliable solution for addressing the disparities arising from differential under-reporting. It underscores the point that such an approach lacks precision, potentially leading to arbitrary and inequitable outcomes.